# Best approximation of a three-variable function by sum of one-variable coordinate functions

### v1

@@ 2.1 Optimal Projection Cycle-Vectors on a Finite Set
$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)\leq\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}|\lambda_{\varepsilon}(x)|=\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon$
@@
$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)=-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon$
@@
$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|\leq-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon\leq 0$
Where the last inequality follows from the fact that the number of summands with a negative sign is $k$, the number of all summands is $m$ and $k\geq\frac{m}{2}$
@@
    Let’s call a plane on $\mathbb{R}^{3}$ an interior plane of the unit cube if it is parallel to one of the faces of the unit cube and passes through an interior point of the cube. If $p$ contains a well-ordered pair of points $x_{i},x_{j}$ lying on the same interior plane, then, $\lambda_{i},\lambda_{j}>0$,
@@
    If $\lambda_{i}<0$, then the point $x_{i}$ belongs to one of the edges of the unit cube not containing any of $(0,0,0)$ and $(1,1,1)$,
@@
    $(0,0,0),(1,1,1)\in p$ and, both has positive weights.
@@
(a) Assume the contrary: Let, $(p,\lambda)=((x_{0},x_{1},...,x_{m}),(\lambda_{0},\lambda_{1},...,\lambda_{m}))\in X_{S,f}$ contain a non well-ordered pair of points $x_{k},x_{j}$ with $\lambda_{k}>0,\lambda_{j}>0$.
Let the coordinates of the points $x_{k},x_{j}$ be $(\xi_{k},\gamma_{k},\zeta_{k})$ and $(\xi_{j},\gamma_{j},\zeta_{j})$ respectively. Since $x_{k},x_{j}$ are not well-ordered, without loss of generality we can assume that $\xi_{k}\geq\xi_{j}$, $\gamma_{k}>\gamma_{j}$ and $\zeta_{k}<\zeta_{j}$.
Denote the points $(\xi_{k},\gamma_{k},\zeta_{j})$ and $(\xi_{j},\gamma_{j},\zeta_{k})$ by $x_{k}^{\prime}$ and $x_{j}^{\prime}$ respectively. Clearly, $x_{k}^{\prime},x_{j}^{\prime}\in S$ and the projection cycle-vector $(p_{0},\lambda_{\varepsilon})\in X_{S}$ with $p_{0}=(x_{k},x_{j},x_{k}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,-\varepsilon,\varepsilon,\varepsilon)$ is well defined for any $\varepsilon>0$. Let, $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$.
Since $\lambda(x_{k})$ and $\lambda(x_{j})$ are positive, when $\varepsilon>0$ is small enough by Lemma $2.1$ we have $0\leq\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$.
On the other hand,
$\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})=\varepsilon(f(x^{\prime}_{k})+f(x^{\prime}_{j})-f(x_{k})-f(x_{j}))>0$
@@
Now since, $\sum\lambda^{\prime}_{i}f(x_{i})>\sum\lambda_{i}f(x_{i})\geq 0$ and $0<\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$, we have
$\frac{\sum\lambda^{\prime}_{i}f(x_{i})}{\sum|\lambda^{\prime}_{i}|}-\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\frac{\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}>0$
Which contradicts to the optimality of $p,\lambda)$.
@@
Without loss of generality we can assume that, $x_{i}$ and $x_{j}$ are lying on the same plane paralel to the $xOy$ plane and thus, we can write, $x_{i}=(\xi_{i},\gamma_{i},\zeta)$ and $x_{j}=(\xi_{j},\gamma_{j},\zeta)$ for some numbers $\xi_{i},\xi_{j},\gamma_{i},\gamma_{j}\in[0,1]$ and some $\zeta\in(0,1)$.
@@ 1. Case 1:
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},0)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},0)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$
  2. Case 2:
@@
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},1)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},1)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,-\varepsilon,-\varepsilon,\varepsilon)$
  3. Case 3:
@@
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{j},\zeta)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{i},\zeta)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,\varepsilon,-\varepsilon,-\varepsilon)$
$x$$z$$y$$x_{i}^{\prime}$$x_{i}$$x_{j}^{\prime}$$x_{j}$
@@
In each case we have chosen $\lambda_{\varepsilon}$ in a way that, at both of the points $x_{i}$ and $x_{j}$ the projection cycle-vectors $(p,\lambda)$ and $(p_{0},\lambda_{\varepsilon})$ have weights with opposite signs. Thus, by the Lemma $2.2$, for $\varepsilon>0$ small enough, we have
@@
Consider the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$. By the definition of the projection cycle each of these planes contains at least one point of $p$ with positive weight. Let these points be $x_{j}=(\xi_{i},\gamma_{j},\zeta_{j}),x_{l}=(\xi_{l},\gamma_{i},\zeta_{l}),x_{k}=(\xi_{k},\gamma_{k},\zeta_{i})$. If any of these positive-weighted points makes a well-ordered pair with $x_{i}$, then the projection cycle-vector $(p,\lambda)$ cannot be optimal by part $(b)$.
$x$$z$$y$$x_{i}$
@@
Assume both $10<\xi<$ 1 and $10<\gamma<$ 1. By the part $(b)$ we have $(x_{i},x_{j})$ is not a well-ordered pair,thus, either $\gamma_{i}<\gamma_{j},\zeta_{i}>\zeta_{j}$ or $\gamma_{i}>\gamma_{j},\zeta_{i}<\zeta_{j}$. Similarly, $(x_{i},x_{l})$ is not a well-ordered pair, thus, either $\xi_{i}<\xi_{l},\zeta_{i}>\zeta_{l}$ or $\xi_{i}>\xi_{l},\zeta_{i}<\zeta_{l}$. It’s easy to verify that in each of the four possible cases, $x_{j},x_{l}$ cannot be a well-ordered pair. By part $(a)$ if $(p,\lambda)$ is optimal then any pair of points $x_{j},x_{l},x_{k}$ should be well-ordered. Thus, $(p,\lambda)$ cannot be optimal, and we get a contradiction. Since, we choose $\xi,\gamma$ arbitrarily from $\xi,\gamma,\zeta$, this contradiction works if at least two of the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$ are inner planes, which happens when $x_{i}$ is not on the edges of the unit cube. Hence, $x_{i}$ belongs to one of the edges of the unit cube.
Now, if, $x_{i}$ and $(0,0,0)$ are on the same edge, we have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$. Assume, $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exist an other point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}>0$. Moreover, since, $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered. But this contradicts the optimality of $(p,\lambda)$, by part $(b)$. Similarly, in the other two cases, we get a contradiction.
@@
Thus, $x_{i}$ belongs to one of the edges of the cube, not containing any of the vertices $(0,0,0)$ and $(1,1,1)$. Let’s name these edges $l_{1},...,l_{6}$ as following
$l_{1}=\{(\xi,0,1),0\leq\xi\leq 1\}$
$l_{2}=\{(1,0,\zeta),0\leq\zeta\leq 1\}$
$l_{3}=\{(1,\gamma,0),0\leq\gamma\leq 1\}$
$l_{4}=\{(\xi,1,0),0\leq\xi\leq 1\}$
$l_{5}=\{(0,1,\zeta),0\leq\zeta\leq 1\}$
$l_{6}=\{(0,\gamma,1),0\leq\gamma\leq 1\}$
@@
If, $x_{i}$ and $(0,0,0)$ are on the same edge, have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$. Assume, $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exist an other point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}<0$. Moreover, since, $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered. But this contradicts the optimality of $(p,\lambda)$, by part $(b)$.
@@
Case 2: $x_{i}$ belongs to one of the edges of the cube, not containing any of the vertices $(0,0,0)$ and $(1,1,1)$. Without loss of generality, we can assume that $x_{i}=(\xi,0,1)$. To make the sum of weights on the plane $z=1$ equal to zero, there should exist a point $x_{j}$ of $p$ with a negative weight on $z=1$. By the part $(c)$ the point $x_{j}$ belongs to at least one of the edges $l_{1}$ or $l_{6}$. Thus, we have two sub-cases:
Case 2.1: $x_{j}=(\xi_{j},0,1)$ for some $0\leq\xi_{j}\leq 1$ Clearly, $\xi_{j}\neq\xi$. We should consider two sub-cases:
@@
Consider the projection cycle $(p,\lambda_{\varepsilon})$, where $p_{0}=(x_{i},x_{j},(\xi,1,1),(\xi_{j},1,1))$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$
@@
By a similar way to the part $(a)$ and part $(b)$, taking $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$ we get a contradiction to the optimality of $(p,\lambda)$.
Case 2.2: $x_{j}=(0,\gamma,1)$ for some $0<\gamma\leq 1$ Consider the plane, $y=\gamma$. To make the sum of weights on the plane $y=\gamma$ equal to zero, there should exist a point $x_{k}$ of $p$ with a positive weight on $y=\gamma$. Let $x_{k}=(\xi_{k},\gamma,\zeta)$.If $\zeta=1$, the points, $x_{j}$ and $x_{k}$ belonging to the inner plane $y=\gamma$ becomes well-ordered and since $x_{j}$ and $x_{k}$ have opposite weights we get a contradiction by part $(b)$. If, $\zeta<1$, the points, $x_{i}$ and $x_{k}$ becomes non well-ordered. Since both of $x_{i}$ and $x_{k}$ have positive weights, we get a contradiction by the part $(a)$.
Since, we get a contradiction in each case, the proof of part $(d)$ is completed.
(e) Consider the point $x_{i}=(\xi_{i},\gamma_{i},\zeta_{i})\in p$ with the minimal coordinates among positive weighted points of $p$. (Since, the positive weighted points are well-ordered pairwise, the point with the minimal sum of coordinates has minimal coordinates.)
If $x_{i}=(0,0,0)$, we are done. Otherwise, since $x_{i}$ is chosen to be minimal, it cannot be any other vertex of the unit cube, and thus there is at least one inner plane passing through $x_{i}$. Without loss of generality, assume this plane has equation $x=\xi_{i}$. Then, on the plane $x=\xi_{i}$ there is another point $x_{j}=(\xi_{i},\gamma_{j},\zeta_{j})\in p$ such that $\lambda_{j}<0$. By part $(b)$ we know that $x_{i}$ and $x_{j}$ cannot be well-ordered and hence, either $\gamma_{j}<\gamma_{i}$ or $\zeta_{j}<\zeta_{i}$ but in the first case on the plane $y=\gamma_{j}$ and in the second case on the plane $z=\zeta_{j}$ there is a positive weighted point to make the sum over that plane equal to zero. But this contradicts to the minimality of $x_{i}$ and hence, $x_{i}=(0,0,0)$.
@@
Let $X^{\prime}_{S}$ be the set of projection cycle-vectors $(p,\lambda)\in X_{S}$, such that the propositions $(a)-(e)$ of the Theorem 2.2 holds for $(p,\lambda)$ and let $X^{\prime}_{S,min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in X^{\prime}_{S}$.
The following lemmas hold.
@@
(See: Lemma 4 and Lemma 5 in [19] or Lemma 3.32 in [11]). ∎
@@
For any projection cycle $(p,\lambda)\in X_{S}$ and for any function $f\in C(\Omega)$, there exist a projection cycle $(p^{0},\lambda^{0})\in X_{S,min}$ such that $p^{0}\subset p$ and the following relations hold
@@
    $\lambda^{0}(x)\lambda(x)>0,\forall x\in p^{0}$
@@
    $\frac{\sum_{x_{j}\in p^{0}}\lambda^{0}_{j}f(x_{j})}{\sum|\lambda^{0}_{j}|}\geq\frac{\sum_{x_{i}\in p}\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
@@
See [19, Theorem 2]. ∎
The Lemma 2.3 implies that, for any function $f$,
$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
@@
$\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$

### v2

@@ 2.1 Optimal Projection Cycle-Vectors on a Finite Set
$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)\leq\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}|\lambda_{\varepsilon}(x)|=\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon.$
@@
$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)=-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon.$
@@
$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|\leq-\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\varepsilon+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\varepsilon+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon\leq 0.$
Where the last inequality follows from the fact that the number of summands with a negative sign is $k$, the number of all summands is $m$, and $k\geq\frac{m}{2}$
@@
    Let us call a plane on $\mathbb{R}^{3}$ an interior plane of the unit cube if it is parallel to one of the faces of the unit cube and passes through an interior point of the cube. If $p$ contains a well-ordered pair of points $x_{i},x_{j}$ lying on the same interior plane, then, $\lambda_{i},\lambda_{j}>0$,
@@
    If $\lambda_{i}<0$, then the point $x_{i}$ belongs to one of the edges of the unit cube that is not adjacent to any of $(0,0,0)$ and $(1,1,1)$,
@@
    Both of the points $(0,0,0)$ and $(1,1,1)$ belong to $p$ and, both have positive weights.
@@
(a) Assume the contrary: Let $(p,\lambda)=((x_{0},x_{1},...,x_{m}),(\lambda_{0},\lambda_{1},...,\lambda_{m}))\in[S]_{f}$ contain a non well-ordered pair of points $x_{k},x_{j}$ with $\lambda_{k}>0,\lambda_{j}>0$.
Let the coordinates of the points $x_{k},x_{j}$ be $(\xi_{k},\gamma_{k},\zeta_{k})$ and $(\xi_{j},\gamma_{j},\zeta_{j})$, respectively. Since $x_{k},x_{j}$ are not well-ordered, without loss of generality we can assume that $\xi_{k}\geq\xi_{j}$, $\gamma_{k}>\gamma_{j}$ and $\zeta_{k}<\zeta_{j}$.
Denote the points $(\xi_{k},\gamma_{k},\zeta_{j})$ and $(\xi_{j},\gamma_{j},\zeta_{k})$ by $x_{k}^{\prime}$ and $x_{j}^{\prime}$, respectively. Clearly, $x_{k}^{\prime},x_{j}^{\prime}\in S$ and the projection cycle-vector $(p_{0},\lambda_{\varepsilon})\in[S]$ with $p_{0}=(x_{k},x_{j},x_{k}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,-\varepsilon,\varepsilon,\varepsilon)$ is well defined for any $\varepsilon>0$. Let, $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$.
Since $\lambda(x_{k})$ and $\lambda(x_{j})$ are positive, when $\varepsilon>0$ is small enough by Lemma 2.2 we have $0\leq\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$. On the other hand,
$\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})=\varepsilon(f(x^{\prime}_{k})+f(x^{\prime}_{j})-f(x_{k})-f(x_{j}))>0,$
@@
Now, since $\sum\lambda^{\prime}_{i}f(x_{i})>\sum\lambda_{i}f(x_{i})\geq 0$ and $0<\sum|\lambda^{\prime}_{i}|\leq\sum|\lambda_{i}|$, we have
$\frac{\sum\lambda^{\prime}_{i}f(x_{i})}{\sum|\lambda^{\prime}_{i}|}-\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\frac{\sum\lambda^{\prime}_{i}f(x_{i})-\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}>0.$
This contradicts the optimality of $(p,\lambda)$.
@@
Without loss of generality, we can assume that $x_{i}$ and $x_{j}$ are lying on the same plane parallel to the $xOy$ plane, and thus we can write $x_{i}=(\xi_{i},\gamma_{i},\zeta)$ and $x_{j}=(\xi_{j},\gamma_{j},\zeta)$ for some numbers $\xi_{i},\xi_{j},\gamma_{i},\gamma_{j}\in[0,1]$ and some $\zeta\in(0,1)$.
@@
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},0)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},0)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$;
  2. 2.
@@
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{i},1)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{j},1)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,-\varepsilon,-\varepsilon,\varepsilon)$;
  3. 3.
@@
    Take $x^{\prime}_{i}=(\xi_{i},\gamma_{j},\zeta)$, $x^{\prime}_{j}=(\xi_{j},\gamma_{i},\zeta)$, $p_{0}=(x_{i},x_{j},x_{i}^{\prime},x_{j}^{\prime})$ and $\lambda_{\varepsilon}=(\varepsilon,\varepsilon,-\varepsilon,-\varepsilon)$.
[…]
@@
In each case we have chosen $\lambda_{\varepsilon}$ in a way that, at both of the points $x_{i}$ and $x_{j}$ the projection cycle-vectors $(p,\lambda)$ and $(p_{0},\lambda_{\varepsilon})$ have weights with opposite signs. Thus, by the Lemma 2.2, for $\varepsilon>0$ small enough, we have
@@
Consider the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$. By the definition of the projection cycle, each of these planes contains at least one point of $p$ with positive weight. Let these points be $x_{j}=(\xi_{i},\gamma_{j},\zeta_{j}),x_{l}=(\xi_{l},\gamma_{i},\zeta_{l}),x_{k}=(\xi_{k},\gamma_{k},\zeta_{i})$. If any of these positive-weighted points makes a well-ordered pair with $x_{i}$, then the projection cycle-vector $(p,\lambda)$ cannot be optimal by part $(b)$.
[…]
@@
Assume both $0<\xi<1$ and $0<\gamma<1$. By the part $(b)$ we have that $(x_{i},x_{j})$ is not a well-ordered pair,thus, either $\gamma_{i}<\gamma_{j},\zeta_{i}>\zeta_{j}$ or $\gamma_{i}>\gamma_{j},\zeta_{i}<\zeta_{j}$. Similarly, $(x_{i},x_{l})$ is not a well-ordered pair, thus, either $\xi_{i}<\xi_{l},\zeta_{i}>\zeta_{l}$ or $\xi_{i}>\xi_{l},\zeta_{i}<\zeta_{l}$. It is easy to verify that in each of the four possible cases, $x_{j},x_{l}$ cannot be a well-ordered pair. By part $(a)$ if $(p,\lambda)$ is optimal, then any pair of points $x_{j},x_{l},x_{k}$ should be well-ordered. Thus, $(p,\lambda)$ cannot be optimal, and we get a contradiction. Since we choose $\xi,\gamma$ arbitrarily from $\xi,\gamma,\zeta$, this contradiction works if at least two of the planes $x=\xi_{i},y=\gamma_{i}$ and $z=\zeta_{i}$ are inner planes, which happens when $x_{i}$ is not on the edges of the unit cube. Hence, $x_{i}$ belongs to one of the edges of the unit cube.
Now, if, $x_{i}$ and $(0,0,0)$ are on the same edge, we have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$. Assume that $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exist another point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}>0$. Moreover, since $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered. But this contradicts the optimality of $(p,\lambda)$, by part $(b)$. Similarly, in the other two cases, we get a contradiction.
@@
Thus, $x_{i}$ belongs to one of the edges of the cube, which does not contain any of the vertices $(0,0,0)$ and $(1,1,1)$. Let’s name these edges $l_{1},...,l_{6}$ as following
$l_{1}=\{(\xi,0,1),0\leq\xi\leq 1\},$
$l_{2}=\{(1,0,\zeta),0\leq\zeta\leq 1\},$
$l_{3}=\{(1,\gamma,0),0\leq\gamma\leq 1\},$
$l_{4}=\{(\xi,1,0),0\leq\xi\leq 1\},$
$l_{5}=\{(0,1,\zeta),0\leq\zeta\leq 1\},$
$l_{6}=\{(0,\gamma,1),0\leq\gamma\leq 1\}.$
@@
If $x_{i}$ and $(0,0,0)$ are on the same edge, have three cases: $x_{i}=(\xi,0,0)$ or $x_{i}=(0,\gamma,0)$ or $x_{i}=(0,0,\zeta)$. Assume $x_{i}=(\xi,0,0)$, then on the inner plane $x=\xi$ there exists another point $x_{j}=(\xi,\gamma_{j},\zeta_{j})\in p$ with $\lambda_{j}<0$. Moreover, since $\gamma_{j},\zeta_{j}\geq 0$, $x_{i}$ and $x_{j}$ are well-ordered. But this contradicts the optimality of $(p,\lambda)$, by part $(b)$.
@@
Case 2: $x_{i}$ belongs to one of the edges of the cube, not containing any of the vertices $(0,0,0)$ and $(1,1,1)$. Without loss of generality, we can assume that $x_{i}=(\xi,0,1)$. To make the sum of weights on the plane $z=1$ equal to zero, there should exist a point $x_{j}$ of $p$ with a negative weight on $z=1$. By part $(c)$ the point $x_{j}$ belongs to at least one of the edges $l_{1}$ or $l_{6}$. Thus, we have two sub-cases:
Case 2.1: $x_{j}=(\xi_{j},0,1)$ for some $0\leq\xi_{j}\leq 1$. Clearly, $\xi_{j}\neq\xi$. We should consider two sub-cases:
@@
Consider the projection cycle $(p,\lambda_{\varepsilon})$, where $p_{0}=(x_{i},x_{j},(\xi,1,1),(\xi_{j},1,1))$ and $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,-\varepsilon)$;
@@
Similarly to part $(a)$ and part $(b)$, taking $(p^{\prime},\lambda^{\prime})=(p,\lambda)+(p_{0},\lambda_{\varepsilon})$ we get a contradiction to the optimality of $(p,\lambda)$.
Case 2.2: $x_{j}=(0,\gamma,1)$ for some $0<\gamma\leq 1$. Consider the plane, $y=\gamma$. To make the sum of weights on the plane $y=\gamma$ equal to zero, there should exist a point $x_{k}$ of $p$ with a positive weight on $y=\gamma$. Let $x_{k}=(\xi_{k},\gamma,\zeta)$. If $\zeta=1$, the points $x_{j}$ and $x_{k}$ belonging to the inner plane $y=\gamma$ becomes well-ordered and since $x_{j}$ and $x_{k}$ have opposite weights we get a contradiction by part $(b)$. If, $\zeta<1$, the points, $x_{i}$ and $x_{k}$ becomes non well-ordered. Since both of $x_{i}$ and $x_{k}$ have positive weights, we get a contradiction by the part $(a)$.
Since we get a contradiction in each case, the proof of part $(d)$ is completed.
(e) Consider the point $x_{i}=(\xi_{i},\gamma_{i},\zeta_{i})\in p$ with minimal coordinates among the positive weighted points of $p$. (Since the positive weighted points are well-ordered pairwise, the point with the minimal sum of coordinates has minimal coordinates.)
If $x_{i}=(0,0,0)$, we are done. Otherwise, since $x_{i}$ is chosen to be minimal, it cannot be any other vertex of the unit cube, and thus there is at least one inner plane passing through $x_{i}$. Without loss of generality, assume that this plane has equation $x=\xi_{i}$. Then, on the plane $x=\xi_{i}$ there is another point $x_{j}=(\xi_{i},\gamma_{j},\zeta_{j})\in p$ such that $\lambda_{j}<0$. By part $(b)$ we know that $x_{i}$ and $x_{j}$ cannot be well-ordered, and hence, either $\gamma_{j}<\gamma_{i}$ or $\zeta_{j}<\zeta_{i}$ but in the first case on the plane $y=\gamma_{j}$ and in the second case on the plane $z=\zeta_{j}$ there is a positive weighted point to make the sum over that plane equal to zero. But this contradicts to the minimality of $x_{i}$ and, hence, $x_{i}=(0,0,0)$.
@@
From now on, $S]^{(a)-(e)}$ will denote the set of projection cycle-vectors $(p,\lambda)\in[S]$, such that the propositions (a)(e)(a)-(e) of the Theorem 2.2 holds for $(p,\lambda)$ and let $S]^{(a)-(e)}_{min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in[S]^{(a)-(e)}$.
The following lemmas will later help us to prove our main results for functions satisfying the inequalities (2.1), instead of the stricter version (2.5).
@@
(See: [31, Lemma 4 and Lemma 5] or [20, Lemma 3.32]). ∎
@@
For any projection cycle $(p,\lambda)\in[S]$ and for any function $f\in C(\Omega)$, there exist a projection cycle $(p^{0},\lambda^{0})\in[S]_{min}$ such that $p^{0}\subset p$ and the following relations hold
@@
    $\lambda^{0}(x)\lambda(x)>0,\ \forall x\in p^{0},$
@@
    $\frac{\sum_{x_{j}\in p^{0}}\lambda^{0}_{j}f(x_{j})}{\sum|\lambda^{0}_{j}|}\geq\frac{\sum_{x_{i}\in p}\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
@@
See [31, Theorem 2]. ∎
Lemma 2.4 implies that, for any function $f$,
$\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
@@
$\max_{(p,\lambda)\in[S]^{(a)-(e)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$